OpenStudy (anonymous):

If each resistor in the circuit drawn inside is 5.0 ohms and the battery is 12.0 V, what does the second ammeter (A_2) read? A. 0.48 A B. 0.24 A C. 0.96 A D. 0.12 A

OpenStudy (anonymous):



OpenStudy (michele_laino):

your circuit is equivalent to this one:  the voltage drop across A and B is \[{V_{AB}} = \frac{{12}}{{5 + 2.5 + 5}} \times 2.5 = ...volts\]